# What Guides Our Choices? Modeling Developers’ Trust and Behavioral Intentions Towards GenAI

### v1

@@ 3.2. Psychometric Analysis of PICSE Framework
Table 4. EFA: Factor loadings and communalities ($h2$)
@@
  • The applied oblique rotation method is promax. Communality values (h2)<0.5 are problematic (Hair, 2009) and are marked in RED. Items loaded well (>0.5) onto their primary factors without cross-loadings (>0.3) onto other factors (Hair et al., 2019); hence their corresponding cells are kept blank.
[…]
3) CFA - Alternate grouping: In the final step, as is best practice (Raykov and Marcoulides, 2011), we conducted CFA to validate the factor structure identified through EFA. The CFA fit indices in Table 5 confirm the EFA-derived four-factor model (RMSEA = 0.048; SRMR = 0.047, CFI = 0.982, TLI = 0.973). Table 4 outlines the factor structure and corresponding item groupings. Factor 1, labeled System/Output quality, includes items S2 through S5 and E3, which relate to the System group (in PICSE) and the style matching of genAI’s outputs. Factor 2, labeled Functional value, encompasses items I3 and P3, reflecting the educational value and practical advantages of using genAI tools. Factor 3, labeled Ease of use, comprises items S1 and C2, addressing the ease of using and integrating genAI in the workflow. Factor 5, labeled Goal maintenance, includes a single item, E4, focusing on genAI’s maintenance of human goals. The reliability and validity assessments further support the robustness of these constructs (see Sec. 4.1).
@@
Table 5. Model Fit Indices - PICSE Psychometric Evaluation
@@
  • $\chi^{2}$ test results were not considered, as the test is affected by deviations from multivariate normality (Schumacker and Lomax, 2004). We still report the values for completeness.
3.3. Model Development
@@
System/Output quality encompasses genAI tools’ presentation, adherence to safe and secure practices (including privacy and security implications of using genAI), and its performance and output quality (consistency and correctness) in relation to the development style or work environment in which it is utilized (S2-S5, E3). Developers often place trust in AI based on its performance and output quality (accuracy and consistency), which serve as proxies for the system’s perceived credibility (Fogg and Tseng, 1999; Wang et al., 2023; Cheng et al., 2023; Yu et al., 2019). Prior work (Wang et al., 2023) evidenced that developers are often wary about the security and privacy implications of using AI tools in their work, which influences the level of trust they place in these tools. Drawing upon these insights, we hypothesize: (H1) System/Output quality of genAI is positively associated with developers’ trust in these tools.
Functional value of a tool refers to the practical benefits and utility it offers users in their work (Sheth et al., 1991). In our context, genAI’s functional value encompasses its educational value and clear advantages relative to work performance (I3, P3). Prior work highlights that developers’ expectations of clear advantages from using AI tools (e.g., increased productivity, improved code quality) contribute to their trust in using these tools (Johnson et al., 2023; Ziegler et al., 2024). Further, AI’s ability to support learning fosters trust in these tools (Wang et al., 2023). Based on these, we posit: (H2) Functional value of genAI is positively associated with developers’ trust in these tools.
Ease of use associated with genAI tools includes the extent to which developers can easily use and integrate genAI into their current workflow (S1, C2). Prior research highlights that a tool’s ease of use (Gefen et al., 2003) and compatibility with existing workflows (Lee and See, 2004; Russo, 2024) contribute to users’ trust. Following this, we hypothesize: (H3) GenAI’s ease of use is positively associated with developers’ trust in these tools.
Goal maintenance is related to the degree to which genAI’s actions and responses align with the developer’s ongoing goals (E4). By its very nature, goals can vary depending on the task and context (Johnson et al., 2023). Therefore, aligning AI behavior with an individual’s immediate goals is crucial in human-AI collaboration scenarios (Wischnewski et al., 2023). In terms of human cognition, this congruence is important for maintaining cognitive flow and reducing cognitive load (Unsworth et al., 2012), which, in turn, fosters trust in systems (van der Werff et al., 2019; Chow and Chan, 2008). Consequently, we propose: (H4) Goal maintenance is positively associated with developers’ trust in genAI tools.
@@
Trust is a key factor in explaining resistance toward automated systems (Wischnewski et al., 2023) and plays an important role in technology adoption (Xiao et al., 2014; Witschey et al., 2015). Multiple studies have correlated an individual’s trust in technology with their intention to use it (Kim et al., 2019; Gefen et al., 2003; Baek and Kim, 2023). In our context, we thus posit: (H5) Trust is positively associated with intentions to use genAI tools.
@@
Motivations behind why someone uses technology (technophilic or task-focused) not only influences their intention to use it but also affects how they engage with its features and functionalities (Venkatesh et al., 2012; O’Brien, 2010). Naturally, individuals motivated by their interest and enjoyment in using and exploring the technology (opposite end of the spectrum from those motivated by task completion) are early adopters of new technology (Burnett et al., 2016). Based on this, we posit: (H6) Motivation to use technology for its own sake is positively associated with intentions to use genAI tools.
@@
Computer self-efficacy refers to an individual’s belief in their ability to engage with and use new technologies to succeed in tasks (Bandura, 1997). It shapes how individuals apply cognitive strategies and the effort and persistence they invest in using new technologies (Compeau and Higgins, 1995), thereby influencing their intention to use them (Venkatesh and Davis, 2000; Li et al., 2024). In line with this, we propose: (H7) Computer self-efficacy is positively associated with intentions to use genAI tools.
@@
Attitude towards risk encompasses an individual’s inclination to take risks in uncertain outcomes (Byrnes et al., 1999). This cognitive facet influences decision-making processes, particularly in contexts involving new or unfamiliar technology (Venkatesh and Davis, 2000). Risk-tolerant individuals (one end of the spectrum) are more inclined to experiment with unproven technology than risk-averse ones (the other end) (Burnett et al., 2016), and show higher intentions to use new tools (Venkatesh et al., 2012; Li et al., 2024). Thus, we posit: (H8) Risk tolerance is positively associated with intentions to use genAI tools.
@@
Information processing style influences how individuals interact with technology when problem-solving: some gather information comprehensively to develop a detailed plan before acting; others gather information selectively, acting on initial promising pieces and acquiring more as needed (Burnett et al., 2016). GenAI systems, by their very interaction paradigm, inherently support the latter by providing immediate responses to queries, allowing users to act quickly on the information received and gather additional details incrementally. Accordingly, we posit: (H9) Selective information processing style is positively associated with intentions to use genAI tools.
@@
Learning style for technology (by process vs. by tinkering) refers to how an individual approaches problem-solving and how they structure their approach to a new technology (Burnett et al., 2016). Some prefer to learn through an organized, step-by-step process, while others prefer to tinker around—exploring and experimenting with new technology or its features (Burnett et al., 2016). Prior work indicates that software, more often than not, is designed to support and encourage tinkering (Carroll and Rosson, 2003), making individuals who prefer this approach more inclined to adopt and use new tools (Venkatesh et al., 2003). Thus, we propose: (H10) Tinkering style is positively associated with intentions to use genAI tools.
@@
Behavioral intention. Successful technology adoption hinges on users’ intention to use it, translating into future usage. Prior work has consistently shown these factors to be positively correlated (Venkatesh et al., 2003, 2012), suggesting that users who intend to use technology are more likely to do so. Accordingly, we hypothesize: (H11) Behavioral intention to use genAI tools is positively associated with the usage of these tools.
3.4. Data Analysis
@@
We used Partial Least Squares-Structural Equation Modeling (PLS-SEM) to test our theoretical model. PLS-SEM is a second-generation multivariate data analysis technique that has gained traction in empirical SE studies investigating complex phenomena (Russo, 2024; Trinkenreich et al., 2023; Russo and Stol, 2021). It allows for simultaneous analysis of relationships among constructs (measured by one or more indicators) and addresses multiple interconnected research queries in one comprehensive analysis. It is particularly suited for exploratory studies due to its flexibility in handling model complexity while accounting for measurement errors in latent variables (Hair et al., 2019). Importantly, PLS-SEM does not require data to meet distributional assumptions. Instead, it uses a bootstrapping approach to determine the statistical significance of path coefficients (i.e., relationships between constructs). The PLS path model is estimated for a large number of random subsamples (usually 5000), generating a bootstrap distribution, which is then used to make statistical inferences (Hair et al., 2019).
@@
We used the SmartPLS (v4.1.0) software (SmartPLS Team, 2024) for PLS-SEM analyses, which comprised two main steps, each involving specific tests and procedures. First, we evaluated the measurement model, empirically assessing the relationships between the latent constructs and their indicators (Sec. 4.1). Next, we evaluated the theoretical (or structural) model (Sec. 4.2), representing the hypotheses presented in Section 3.3.
@@
The appropriate sample size was determined by conducting power analysis using the G*Power tool  (Faul et al., 2009). We performed an F-test with multiple linear regression, setting a medium effect size (0.25), a significance level of 0.05, and a power of 0.95. The maximum number of predictors in our model is seven (six theoretical constructs and one control variable to Behavioral Intention) (see Fig. 1). The calculation indicated a minimum sample size of 95; our final sample size of 238 exceeded it considerably.
4. Results
@@
In this section, we report the evaluation of the measurement model (Sec. 4.1), followed by the evaluation of the structural model (Sec. 4.2). We adhered to the evaluation protocols outlined in prior studies (Russo and Stol, 2021; Hair et al., 2019). The analysis was performed using the survey data, which met the assumptions for factor analysis (Hair et al., 2019): significant Bartlett’s test of sphericity on all constructs ($\chi^{2}$(496)=4474.58, p $<$ .001) and adequate KMO measure of sampling adequacy (0.901), well above the recommended threshold (0.60) (Howard, 2016).
4.1. Measurement Model Evaluation
@@
Our model evaluates several theoretical constructs that are not directly observable (e.g., Trust, Behavioral Intention). These constructs are modeled as latent variables, each measured by a set of indicators or manifest variables (see Fig. 1). The first step in evaluating a structural equation model is to ensure the soundness of the measurement of these latent variables, a process referred to as evaluating the ‘measurement model’ (Hair et al., 2019). We performed a series of tests to validate the measurement model (Russo and Stol, 2021), detailed as follows:
1) Convergent validity examines how a measure correlates with alternate measures of the same construct, focusing on the correlations between indicators (questions) and their corresponding construct. This evaluation assesses whether respondents interpret the questions as intended by the question designers (Kock, 2014). Our theoretical model comprises latent constructs that are reflectively measured, meaning the changes in the construct should be reflected in changes in the indicators (Russo and Stol, 2021). Consequently, these indicators should exhibit a significant proportion of shared variance by converging on their respective constructs (Hair et al., 2019). We assessed convergent validity using Average Variance Extracted (AVE) and indicator reliability through outer loadings (Hair et al., 2019).
AVE represents a construct’s communality, indicating the shared variance among its indicators, and should exceed 0.5 (Hair et al., 2019). AVE values for all latent constructs in our model surpassed this threshold (see Table 6). Regarding outer loadings, values above 0.708 are considered sufficient, while values above 0.60 are sufficient for exploratory studies (Hair et al., 2019). We removed variables that did not sufficiently reflect changes in the latent construct (SE3 from computer self-efficacy and IP3 from selective information processing). [^2] After removing SE3 and IP3, the AVE values for computer self-efficacy (now with 3 indicators) and selective information processing (now with 2 indicators) increased from 0.627 to 0.736 and 0.609 to 0.741, respectively. Thus, all indicators in our model exceeded the threshold, ranging between 0.615 and 0.954 (see Fig. 1).
Table 6. Internal consistency reliability and convergent validity
@@
  • Cronbach’s $\alpha$ tends to underestimate reliability, whereas composite reliability (CR: $\rho_{c}$) tends to overestimate it. The true reliability typically lies between these two estimates and is effectively captured by CR($\rho_{a}$) (Russo and Stol, 2021).
@@
2) Internal consistency reliability seeks to confirm that the indicators are consistent with one another and that they consistently and reliably measure the same construct. To assess this, we performed both Cronbach’s $\alpha$ and Composite Reliability (CR: $\rho_{a},\rho_{c}$) tests (Russo and Stol, 2021). The desirable range for these values is between 0.7 and 0.9 (Hair et al., 2019). As presented in Table 6, all values corresponding to our model constructs fall within the acceptable range, confirming that the constructs and their indicators meet the reliability criteria.
@@
3) Discriminant validity assesses the distinctiveness of each construct in relation to the others. Our model includes 10 latent variables (Table 6). A primary method for assessing discriminant validity is the Heterotrait-Monotrait (HTMT) ratio of correlations (Henseler et al., 2015). Discriminant validity may be considered problematic if the HTMT ratio $>0.9$, with a more conservative cut-off at $0.85$ (Hair et al., 2019). In our case, the HTMT ratios between the latent constructs ranged from 0.064 to 0.791, all below the threshold. We report the HTMT ratios in the supplemental (sup, [n. d.]), along with the cross-loadings of the indicators, and the Fornell-Larcker criterion values for the sake of completeness.
@@
4) Collinearity assessment is conducted to evaluate the correlation between predictor variables, ensuring they are independent to avoid potential bias in the model path estimations. We assessed collinearity using the Variance Inflation Factor (VIF). In our model, all VIF values are below 2.1, well below the accepted cut-off value of 5 (Hair et al., 2019).
@@
Figure 1. PLS-SEM Model: Solid lines indicate item loadings and path coefficients (p $<$ 0.05); dashed lines represent non-significant paths. Reverse-coded items are suffixed with ‘-R’ (e.g., SE2-R). Latent constructs are depicted as circles and adjusted $R^{2}$ (Adj. $R^{2}$) values are reported for endogenous constructs.
4.2. Structural Model Evaluation

### v2

@@ III-B Psychometric Analysis of PICSE Framework
TABLE IV: EFA: Factor loadings and communalities ($h2$)
@@
  • The applied oblique rotation method is promax. Communality values (h2)<0.5 are problematic [64] and are marked in RED. Items loaded well (>0.5) onto their primary factors without cross-loadings (>0.3) onto other factors [52]; hence their corresponding cells are kept blank.
3) CFA - Alternate grouping: In the final step, as is best practice [58], we conducted CFA to validate the factor structure identified through EFA. The CFA fit indices in Table V confirm the EFA-derived four-factor model (RMSEA = 0.048; SRMR = 0.047, CFI = 0.982, TLI = 0.973). Table IV outlines the factor structure and corresponding item groupings. Factor 1, labeled System/Output quality, includes items S2 through S5 and E3, which relate to the System group (in PICSE) and the style matching of genAI’s outputs. Factor 2, labeled Functional value, encompasses items I3 and P3, reflecting the educational value and practical advantages of using genAI tools. Factor 3, labeled Ease of use, comprises items S1 and C2, addressing the ease of using and integrating genAI in the workflow. Factor 5, labeled Goal maintenance, includes a single item, E4, focusing on genAI’s maintenance of human goals. The reliability and validity assessments further support the robustness of these constructs (see Sec. IV-A).
@@
TABLE V: Model Fit Indices - PICSE Psychometric Evaluation
@@
  • $\chi^{2}$ test results were not considered, as the test is affected by deviations from multivariate normality [65]. We still report the values for completeness.
III-C Model Development
@@
System/Output quality encompasses genAI tools’ presentation, adherence to safe and secure practices (including privacy and security implications of using genAI), and its performance and output quality (consistency and correctness) in relation to the development style or work environment in which it is utilized (S2-S5, E3). Developers often place trust in AI based on its performance and output quality (accuracy and consistency), which serve as proxies for the system’s perceived credibility [66, 14, 15, 67]. Prior work [14] evidenced that developers are often wary about the security and privacy implications of using AI tools in their work, which influences the level of trust they place in these tools. Drawing upon these insights, we hypothesize: (H1) System/Output quality of genAI is positively associated with developers’ trust in these tools.
Functional value of a tool refers to the practical benefits and utility it offers users in their work [68]. In our context, genAI’s functional value encompasses its educational value and clear advantages relative to work performance (I3, P3). Prior work highlights that developers’ expectations of clear advantages from using AI tools (e.g., increased productivity, improved code quality) contribute to their trust in using these tools [16, 69]. Further, AI’s ability to support learning fosters trust in these tools [14]. Based on these, we posit: (H2) Functional value of genAI is positively associated with developers’ trust in these tools.
Ease of use associated with genAI tools includes the extent to which developers can easily use and integrate genAI into their current workflow (S1, C2). Prior research highlights that a tool’s ease of use [70] and compatibility with existing workflows [10, 24] contribute to users’ trust. Following this, we hypothesize: (H3) GenAI’s ease of use is positively associated with developers’ trust in these tools.
Goal maintenance is related to the degree to which genAI’s actions and responses align with the developer’s ongoing goals (E4). By its very nature, goals can vary depending on the task and context [16]. Therefore, aligning AI behavior with an individual’s immediate goals is crucial in human-AI collaboration scenarios [34]. In terms of human cognition, this congruence is important for maintaining cognitive flow and reducing cognitive load [71], which, in turn, fosters trust in systems [72, 73]. Consequently, we propose: (H4) Goal maintenance is positively associated with developers’ trust in genAI tools.
@@
Trust is a key factor in explaining resistance toward automated systems [34] and plays an important role in technology adoption [74, 75]. Multiple studies have correlated an individual’s trust in technology with their intention to use it [76, 70, 77]. In our context, we thus posit: (H5) Trust is positively associated with intentions to use genAI tools.
@@
Motivations behind why someone uses technology (technophilic or task-focused) not only influences their intention to use it but also affects how they engage with its features and functionalities [22, 78]. Naturally, individuals motivated by their interest and enjoyment in using and exploring the technology (opposite end of the spectrum from those motivated by task completion) are early adopters of new technology [26]. Based on this, we posit: (H6) Motivation to use technology for its own sake is positively associated with intentions to use genAI tools.
@@
Computer self-efficacy refers to an individual’s belief in their ability to engage with and use new technologies to succeed in tasks [79]. It shapes how individuals apply cognitive strategies and the effort and persistence they invest in using new technologies [80], thereby influencing their intention to use them [81, 82]. In line with this, we propose: (H7) Computer self-efficacy is positively associated with intentions to use genAI tools.
@@
Attitude towards risk encompasses an individual’s inclination to take risks in uncertain outcomes [83]. This cognitive facet influences decision-making processes, particularly in contexts involving new or unfamiliar technology [81]. Risk-tolerant individuals (one end of the spectrum) are more inclined to experiment with unproven technology than risk-averse ones (the other end) [26], and show higher intentions to use new tools [22, 82]. Thus, we posit: (H8) Risk tolerance is positively associated with intentions to use genAI tools.
@@
Information processing style influences how individuals interact with technology when problem-solving: some gather information comprehensively to develop a detailed plan before acting; others gather information selectively, acting on initial promising pieces and acquiring more as needed [26]. GenAI systems, by their very interaction paradigm, inherently support the latter by providing immediate responses to queries, allowing users to act quickly on the information received and gather additional details incrementally. Accordingly, we posit: (H9) Selective information processing style is positively associated with intentions to use genAI tools.
@@
Learning style for technology (by process vs. by tinkering) refers to how an individual approaches problem-solving and how they structure their approach to a new technology [26]. Some prefer to learn through an organized, step-by-step process, while others prefer to tinker around—exploring and experimenting with new technology or its features [26]. Prior work indicates that software, more often than not, is designed to support and encourage tinkering [84], making individuals who prefer this approach more inclined to adopt and use new tools [21]. Thus, we propose: (H10) Tinkering style is positively associated with intentions to use genAI tools.
@@
Behavioral intention. Successful technology adoption hinges on users’ intention to use it, translating into future usage. Prior work has consistently shown these factors to be positively correlated [21, 22], suggesting that users who intend to use technology are more likely to do so. Accordingly, we hypothesize: (H11) Behavioral intention to use genAI tools is positively associated with the usage of these tools.
III-D Data Analysis
@@
We used Partial Least Squares-Structural Equation Modeling (PLS-SEM) to test our theoretical model. PLS-SEM is a second-generation multivariate data analysis technique that has gained traction in empirical SE studies investigating complex phenomena [24, 55, 85]. It allows for simultaneous analysis of relationships among constructs (measured by one or more indicators) and addresses multiple interconnected research queries in one comprehensive analysis. It is particularly suited for exploratory studies due to its flexibility in handling model complexity while accounting for measurement errors in latent variables [52]. Importantly, PLS-SEM does not require data to meet distributional assumptions. Instead, it uses a bootstrapping approach to determine the statistical significance of path coefficients (i.e., relationships between constructs). The PLS path model is estimated for a large number of random subsamples (usually 5000), generating a bootstrap distribution, which is then used to make statistical inferences [52].
@@
We used the SmartPLS (v4.1.0) software [86] for PLS-SEM analyses, which comprised two main steps, each involving specific tests and procedures. First, we evaluated the measurement model, empirically assessing the relationships between the latent constructs and their indicators (Sec. IV-A). Next, we evaluated the theoretical (or structural) model (Sec. IV-B), representing the hypotheses presented in Section III-C.
@@
The appropriate sample size was determined by conducting power analysis using the G*Power tool  [87]. We performed an F-test with multiple linear regression, setting a medium effect size (0.25), a significance level of 0.05, and a power of 0.95. The maximum number of predictors in our model is seven (six theoretical constructs and one control variable to Behavioral Intention) (see Fig. 1). The calculation indicated a minimum sample size of 95; our final sample size of 238 exceeded it considerably.
IV Results
@@
In this section, we report the evaluation of the measurement model (Sec. IV-A), followed by the evaluation of the structural model (Sec. IV-B). We adhered to the evaluation protocols outlined in prior studies [85, 52]. The analysis was performed using the survey data, which met the assumptions for factor analysis [52]: significant Bartlett’s test of sphericity on all constructs ($\chi^{2}$(496)=4474.58, p $<$ .001) and adequate KMO measure of sampling adequacy (0.901), well above the recommended threshold (0.60) [60].
IV-A Measurement Model Evaluation
@@
Our model evaluates several theoretical constructs that are not directly observable (e.g., Trust, Behavioral Intention). These constructs are modeled as latent variables, each measured by a set of indicators or manifest variables (see Fig. 1). The first step in evaluating a structural equation model is to ensure the soundness of the measurement of these latent variables, a process referred to as evaluating the ‘measurement model’ [52]. We performed a series of tests to validate the measurement model [85], detailed as follows:
TABLE VI: Internal Consistency Reliability and Convergent Validity
@@
  • Cronbach’s $\alpha$ tends to underestimate reliability, whereas composite reliability (CR: $\rho_{c}$) tends to overestimate it. The true reliability typically lies between these two estimates and is effectively captured by CR($\rho_{a}$) [85].
1) Convergent validity examines how a measure correlates with alternate measures of the same construct, focusing on the correlations between indicators (questions) and their corresponding construct. This evaluation assesses whether respondents interpret the questions as intended by the question designers [88]. Our theoretical model comprises latent constructs that are reflectively measured, meaning the changes in the construct should be reflected in changes in the indicators [85]. Consequently, these indicators should exhibit a significant proportion of shared variance by converging on their respective constructs [52]. We assessed convergent validity using Average Variance Extracted (AVE) and indicator reliability through outer loadings [52].
AVE represents a construct’s communality, indicating the shared variance among its indicators, and should exceed 0.5 [52]. AVE values for all latent constructs in our model surpassed this threshold (see Table VI). Regarding outer loadings, values above 0.708 are considered sufficient, while values above 0.60 are sufficient for exploratory studies [52]. We removed variables that did not sufficiently reflect changes in the latent construct (SE3 from computer self-efficacy and IP3 from selective information processing). [^2] After removing SE3 and IP3, the AVE values for computer self-efficacy (now with 3 indicators) and selective information processing (now with 2 indicators) increased from 0.627 to 0.736 and 0.609 to 0.741, respectively. Thus, all indicators in our model exceeded the threshold, ranging between 0.615 and 0.954 (see Fig. 1).
@@
2) Internal consistency reliability seeks to confirm that the indicators are consistent with one another and that they consistently and reliably measure the same construct. To assess this, we performed both Cronbach’s $\alpha$ and Composite Reliability (CR: $\rho_{a},\rho_{c}$) tests [85]. The desirable range for these values is between 0.7 and 0.9 [52]. As presented in Table VI, all values corresponding to our model constructs fall within the acceptable range, confirming that the constructs and their indicators meet the reliability criteria.
@@
3) Discriminant validity assesses the distinctiveness of each construct in relation to the others. Our model includes 10 latent variables (Table VI). A primary method for assessing discriminant validity is the Heterotrait-Monotrait (HTMT) ratio of correlations [89]. Discriminant validity may be considered problematic if the HTMT ratio $>0.9$, with a more conservative cut-off at $0.85$ [52]. In our case, the HTMT ratios between the latent constructs ranged from 0.064 to 0.791, all below the threshold. We report the HTMT ratios in the supplemental [53], along with the cross-loadings of the indicators, and the Fornell-Larcker criterion values for the sake of completeness.
@@
4) Collinearity assessment is conducted to evaluate the correlation between predictor variables, ensuring they are independent to avoid potential bias in the model path estimations. We assessed collinearity using the Variance Inflation Factor (VIF). In our model, all VIF values are below 2.1, well below the accepted cut-off value of 5 [52].
@@
Figure 1: PLS-SEM Model: Solid lines indicate item loadings and path coefficients (p $<$ 0.05); dashed lines represent non-significant paths. Reverse-coded items are suffixed with ‘-R’ (e.g., SE2-R). Latent constructs are depicted as circles and adjusted $R^{2}$ (Adj. $R^{2}$) values are reported for endogenous constructs.
IV-B Structural Model Evaluation